hi I'm doing algebra and I'm on evaluating expressions and I don't get the problem it says 0.4r when R=6
Answer by stanbon(75887)

0.4r when R=6
----
Ans: 0.4R = 0.4*6 = (4/10)*6 = 24/10 = 2.4
===============================
Cheers,
Stan H.